A stone thrown horizontally from a height of 6.02 m hits the ground at a distance of 13.50 m. Calculate the initial speed of the stone. Neglect air resistance.

Calculate the speed of the stone as it hits the ground. Neglect air resistance.
1
Expert's answer
2012-11-06T09:20:18-0500
let us find time of falling. height is related with time as
h=gt^2/2
Hence
t=sqrt(2h/g) = sqrt(2*6.02/9.8) = 1.1 s
The horizontal (initial) speed is
13.5/1.1 =12.3 m/s